OpenStudy (anonymous):

Can anyone help with this question? Derive the equation of the parabola with a focus at (-7, 5) and a directrix of y = -11 a. 1/32(x+7)^2 - 3 b. -1/32(x + 7)^2 -3 c. -1/32(x-7)^2 - 3 d. 1/32(x-7)^2 -3

OpenStudy (anonymous):

what would be the formula I would use?

hero (hero):

@annamasreya, are you still here?

OpenStudy (anonymous):

Yes Hero. Hi! I think this is a graphing functions question but don't know how to approach it. I am an online/homeschooled

hero (hero):

In general if you are given a focus point \((\color\green{x_1}, \color\red{y_1})\) and a directrix point \((\color\green{x_2}, \color\red{y_2})\), you can insert both points in to the following formula: \((x - \color\green{x_1})^2 + (y - \color\red{y_1})^2 = (x - \color\green{x_2})^2 + (y - \color\red{y_2})^2\)

hero (hero):

Afterwards, you would simplify the equation accordingly. In this case, we are given: Focus = (-7,5) Directrix = (x, -11)

OpenStudy (anonymous):

ok I am with you so far

OpenStudy (anonymous):

what would we put in for the x on the second set of coordinates? something that would work in the formula?

hero (hero):

And if we insert both points in to the equation we have: \((x - \color\green{(-7)})^2 + (y - \color\red{(5)})^2 = (x - \color\green{(x)})^2 + (y - \color\red{(-11)})^2\)

OpenStudy (anonymous):

ok and just factor them out to get a quadratic statement?

hero (hero):

Basically, it simplifies to \((x + 7)^2 + (x + 5)^2 = (0)^2 + (y + 11)^2\) or just \((x + 7)^2 + (y + 5)^2 = (y + 11)^2\)

hero (hero):

At this point what you want to do is expand \((y + 5)^2\) and \((y + 11)^2\)

OpenStudy (anonymous):

it's asking to get the f(x)? I would get y^2 +10y +25 and y^2 +22y + 121?

hero (hero):

Okay, so what you have at this point is: \((x + 7)^2 + y^2 + 10y + 25 = y^2 + 22y + 121\)

hero (hero):

Now subtract \(y^2 + 10y + 25\) from both sides.

OpenStudy (anonymous):

I think I would end up with (x+7)^2 = 12y +121?

OpenStudy (anonymous):

or do I factor out the (x+7)^2?

hero (hero):

You forgot to subtract 25 from 121

OpenStudy (anonymous):

oops yep so 12y +96

hero (hero):

So you now have \((x + 7)^2 = 12y + 96\) At this point what you want to do is isolate \(y\) as this was the ultimate goal.

hero (hero):

Actually hang on a minute...

OpenStudy (anonymous):

ok

hero (hero):

We should have gotten: \((x + 7)^2 + (y - 5)^2 = (y + 11)^2\) Which expands to: \((x + 7)^2 + y^2 - 10y + 25 = y^2 + 22y + 121\)

hero (hero):

Now subtract \(y^2 - 10y + 25\) from both sides.

OpenStudy (anonymous):

so (x+7)^2 = 12y + 96? I think?

hero (hero):

Had you subtracted correctly, you would have gotten \((x + 7)^2 = 32y + 96\)

hero (hero):

From there, you can solve for y

OpenStudy (anonymous):

so I missed adding the 10y's together. I see

OpenStudy (anonymous):

that would have got my 32y?

hero (hero):

Continue solving for y

OpenStudy (anonymous):

so subtract 96 from both sides then we get (x+7)^2 -96 over 32 = y?

hero (hero):

Yes, but it can be simplifed further

OpenStudy (anonymous):

hmmm. (x+7)^2 -3 over 32????? : )!?

hero (hero):

\(\dfrac{(x + 7)^2}{32} - 3 = y\)

hero (hero):

That result should agree with one of your answer choices.

OpenStudy (anonymous):

Yes! Thank you soooooo much! I feel like we were on a quest and we just slayed the dragon!!!!

OpenStudy (anonymous):

Thank you and I hope you are a teacher because you can really help people!

hero (hero):

Which answer choice is correct?

OpenStudy (anonymous):

it's A.

hero (hero):

Yes, A is correct.

OpenStudy (anonymous):

Thank you for the help! Have a good one!